When calcium carbonate is added to hydrochloric acid the products are calcium chloride, carbon dioxide, and water.

How many grams of calcium chloride will be produced when 29.0 g of calcium carbonate are combined with 11.0 g of hydrochloric acid?

Which reactant is in excess and how many grams of this reactant will remain after the reaction is complete?
Answer #1

CaCO3 + 2HCl -------------> CaCl2 + CO2 + H2O

100.1 g CaCO3 reacts with 2 x 36.5 g HCl

X g CaCO3 reacts with 11.0 g HCl

X = 11 x 100.1 / 2 x 36.5 = 15.08 CaCO3

but we have 29.0 g CaCO3 so CaCO3 is exess reagent

exess CaCO3 left 29.0 - 15.08 = 13.92 g

2 x 36.5 g HCl gives 110.98 g CaCl2

11.0 g HCl gives 11.0 x 110.98 / 73 = 16.72 g

mass of CaCl2 formed = 16.72 g